DETECTING HYDROGEN ATOMS IN SOLID PARAHYDROGEN USING FTIR SPECTROSCOPY

\underline{DAVID T. ANDERSON} and MAHMUT RUZI, Department of Chemistry, University of Wyoming, Laramie, WY 82071-3838.

Our group is currently studying the 193.3 nm photochemistry of a number of precursor molecules isolated in solid parahydrogen (pH_{2}) using high-resolution FTIR spectroscopy. In photochemical studies of formic acid^{a} and ammonia, right after the photolysis laser is turned off, we observe metastable satellite peaks in close proximity to the strong rovibrational transition out of the ground state of the respective monomer photoproducts H_{2}O and NH_{3}. In both cases, there are two satellite peaks near the R(0) monomer transition. We have assigned these satellite features to H-H_{2}O and H-NH_{3} radical clusters that form via reactions of the photoproduct with the pH_{2} host. We will present IR spectroscopic studies of these two H-atom cluster systems that we hope to use in future photochemical studies to measure the concentration of H-atoms in our sample using FTIR spectroscopy.

aK. A. Kufeld, W. R. Wonderly, L. O. Paulson, S. C. Kettwich, and D. T. Anderson, J. Phys. Chem. Lett. 3, 342-347 (2012).